OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (anonymous):

A. y = −4x + 20 B. y = 4x + 20 C. y = −4x – 35 D. y = 4x – 35 E. y = -4x + 4

OpenStudy (anonymous):

d

OpenStudy (anonymous):

jk idk